Suppose 500 mL of 0.200 M nitrous acid , which has a pKa of 3.347, solution is titrated with 75.0 mL of a 1.00 M NaOH. What is the final pH? The error interval is +/- 0.2 pH units.
Answer #1

its a buffer solution

Millimnoles of Nitrous acid taken= 100

Millimoles of NaOH = 75

Millimiles of Salt formed   = 100-75 = 75

Millimoles of Nitrous acid left unreacted = 25

pH = pKa + log(salt/acid) = 3.347 + log(75/25)

                                       = 3.347 + 0.4771 =3.8241